# Text-to-State Mapping for Non-Resolution Reasoning: A Formal Framework for Collapse Deferment in LLM Inference

### v1

@@ 1.2 Non-Resolution Reasoning
Non-Resolution Reasoning (NRR) [1] offers an alternative paradigm that treats ambiguity preservation as a valid computational mode. NRR introduced the state space $\mathcal{S}=\{(v_{i},c_{i},w_{i})\}$ where multiple interpretations coexist with independent activation weights, and demonstrated that prompt-based “NRR-lite” implementations maintain high entropy ($H=0.63$) at ambiguous turns while baseline models collapse ($H=0.10$).
@@ 1.3 Contributions
    Experimental Validation: We evaluate on 68 ambiguous sentences using three LLMs (GPT-4, Gemini, Claude), demonstrating consistent entropy preservation ($H=1.087$) across ambiguity categories.
@@ 2.4 Relation to NRR Framework
This paper builds directly on the NRR framework [1]. Paper 1 established the state space formalism and demonstrated entropy preservation through prompt engineering. NRR defines eight operators ($\sigma,\alpha,\rho,\iota,\delta,\tau,\kappa,\pi$) for manipulating non-collapsing states (see Appendix D for summary).
@@ 5.1 Research Questions
  • RQ4: How consistent are results across different LLMs (GPT-4, Gemini, Claude)?
@@ 5.5 Results: LLM-Based Extraction
To address the limitations of rule-based extraction, we evaluated LLM-based interpretation extraction using three models: GPT-4, Gemini, and Claude. Each model was prompted to enumerate all possible interpretations for 28 English sentences (8 epistemic + 10 lexical + 10 structural).
Category GPT-4 Gemini Claude Mean
@@
Figure 2: LLM-based extraction results across three models (GPT-4, Gemini, Claude) for English sentences. The dashed line indicates $H_{\max}=1.0$ for binary interpretations. All models achieve $H>0$ across all categories, with epistemic showing the highest entropy due to multiple interpretive layers.
@@ 5.6 Combined Results
Overall 68 2.24 1.087 Hybrid
Table 4: Combined results using hybrid extraction. All categories achieve $|S|\geq 2$ and $H>0$, confirming Theorem 1.
Figure 3: Combined experimental results. (A) State entropy $H(S)$ by category, showing rule-based extraction (teal) achieves $H\approx 1.0$ for adversative and hedging, while LLM-based extraction (coral) captures epistemic, lexical, and structural ambiguity. (B) Number of interpretations extracted per category. (C) Overall comparison: baseline systems collapse to $H=0$, while $\phi$ preserves $H=1.087$ across all 68 sentences.
5.7 Baseline Comparison
For comparison, a collapse-based system that commits to a single interpretation yields $H_{\text{baseline}}=0$ for all inputs. Note that this baseline entropy is a theoretical value derived from the architectural constraint that standard LLMs maintain a single active interpretation during inference; we do not measure $H$ empirically for baseline systems, as they do not expose internal interpretive states. The improvement from $\phi$ is:
$\Delta H=H(\phi(T))-H_{\text{baseline}}=1.087-0=+1.087\text{ bits}$ (15)
This represents substantial information preservation. For epistemic sentences specifically, $\Delta H=1.673$ bits—exceeding the theoretical maximum for binary states ($H_{\max}=1.0$) because LLMs extract 3+ interpretations on average.
Answering the Research Questions.
(RQ1) Yes, $\phi$ preserves interpretive multiplicity: mean $|S|=2.24$ and $H=1.087$ across all categories, compared to $|S|=1$ and $H=0$ for baselines.
@@ 5.7 Baseline Comparison
(RQ4) All three LLMs (GPT-4, Gemini, Claude) achieve $H>0$ for 100% of English test sentences. Claude extracts the most interpretations (mean $|S|=3.75$ for epistemic), while Gemini is more conservative ($|S|=3.00$). Despite variation in $|S|$, all models successfully preserve interpretive multiplicity.

### v2

# NRR-Phi: Text-to-State Mapping for Ambiguity Preservation in LLM Inference

@@ 1.2 Non-Resolution Reasoning
Non-Resolution Reasoning (NRR) [1] offers an alternative paradigm that treats ambiguity preservation as a valid computational mode. NRR introduced the state space $\mathcal{S}=\{(v_{i},c_{i},w_{i})\}$ where multiple interpretations coexist with independent activation weights, and demonstrated that “NRR-lite” implementations (k=2k{=}2 separate embeddings with context gate) maintain high entropy ($H=0.91$ bits, near-maximum $1.0$) at ambiguous turns while baseline models collapse ($H=0.15$ bits).
@@ 1.3 Contributions
    Experimental Validation: We evaluate on 68 ambiguous sentences using three LLMs (ChatGPT, Gemini, Claude), demonstrating consistent entropy preservation ($H=1.087$) across ambiguity categories.
@@ 2.4 Relation to NRR Framework
This paper builds directly on NRR-Core [1]. That work established the state space formalism and demonstrated entropy preservation through prompt engineering. NRR defines eight operators ($\sigma,\alpha,\rho,\iota,\delta,\tau,\kappa,\pi$) for manipulating non-collapsing states (see Appendix D for summary).
@@ 5.1 Research Questions
  • RQ4: How consistent are results across different LLMs (ChatGPT, Gemini, Claude)?
@@ 5.3 Metrics
Note that $H_{\max}=\log_{2}|S|$ depends on the number of interpretations; $H_{\max}=1.0$ bit applies only to binary ($|S|=2$) states. States with $|S|>2$ can achieve $H>1.0$ bits.
@@ 5.5 Results: LLM-Based Extraction
To address the limitations of rule-based extraction, we evaluated LLM-based interpretation extraction using three models: ChatGPT, Gemini, and Claude. [^1] All experiments used the free tier of each service’s web interface (January 2026). Specific version numbers are not displayed in free-tier interfaces; based on publicly documented model availability, the models were approximately: ChatGPT (GPT-5.2), Claude (Claude 3.7 Sonnet or later), and Gemini (Gemini 2.5 Flash). Each model was prompted to enumerate all possible interpretations for 28 English sentences (8 epistemic + 10 lexical + 10 structural).
Category ChatGPT Gemini Claude Mean
@@
Figure 2: LLM-based extraction results across three models (ChatGPT, Gemini, Claude) for English sentences. The dashed line indicates $H_{\max}=1.0$ for binary interpretations. All models achieve $H>0$ across all categories, with epistemic showing the highest entropy due to multiple interpretive layers.
@@ 5.6 Combined Results
Overall 68 2.34 1.087 Hybrid
Table 4: Combined results using hybrid extraction. All categories achieve $H>0$, with overall mean entropy $H=1.087$ bits.
Figure 3: Summary of experimental results. (A) State entropy by category and extraction method. Rule-based extraction achieves $H\approx 1.0$ for adversative and hedging; LLM-based extraction achieves $H>0$ for all categories. (B) Mean number of interpretations extracted per category. Epistemic shows highest $|S|$ due to multiple interpretive layers. (C) Overall comparison: baseline (single interpretation, $H=0$) vs. rule-based ($H=1.019$) vs. LLM-based ($H=1.183$) vs. combined ($H=1.087$).
5.7 Research Question Answers
(RQ1) Yes. The mapping $\phi$ preserves interpretive multiplicity: mean $H=1.087$ bits across all 68 sentences, compared to $H=0$ for collapse-based baselines.
@@ 5.7 Research Question Answers
(RQ4) All three LLMs (ChatGPT, Gemini, Claude) achieve $H>0$ for 100% of English test sentences. Claude extracts the most interpretations (mean $|S|=3.75$ for epistemic), while Gemini is more conservative ($|S|=3.00$). Despite variation in $|S|$, all models successfully preserve interpretive multiplicity.
@@ 5.8 Limitations
Note on LLM role: LLMs serve as convenient sources of human-like interpretation enumeration; the theoretical claims of this paper concern the structure of the resulting states, not the capabilities of the underlying models. The $\phi$ mapping framework is model-agnostic and could be instantiated with any interpretation source.